OpenStudy (blazeryder):

If 2.0 J of work is done in raising a 180 g apple, how far is it lifted?

OpenStudy (blazeryder):

I have 2 formulas: \[W=F \times D\] and \[W _{net}=F _{net} \times d \times \cos \theta\]

OpenStudy (osprey):

work done = potential energy gained = mgh. work = 2 joules m=180grams , g = 10m/s/s, .... h = 2/(180x0.001x10)metres. The mass is expressed in kilograms. http://perendis.webs.com

OpenStudy (jiteshmeghwal9):

*work done = -potential energy gained @osprey

OpenStudy (abhilashkk0309):

Work done = P.E gained , So mgh = W ; 180*9.81*h = 2